OpenStudy (anonymous):

Find the limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}(6x+1)/(3x+8)\]

OpenStudy (anonymous):

do it with your eyeballs degree of numerator is the same as the degree of the denominator, (they are both one) so take the ratio of the leading coefficients

OpenStudy (anonymous):

in your case it is \(\frac{6}{3}=2\)

OpenStudy (anonymous):

that easy huh? what about something with different degrees? like \[\lim_{x \rightarrow \infty}(3x ^{2}+1)/x\]

OpenStudy (anonymous):

then you have to compare degrees: higher numerator, limit is infinite ( or doesn't exist) higher denominator, limit is zero

OpenStudy (anonymous):

here's the rule: if the degree on the top = degree on the bottom, then the limit is the ratio of the coefficients of the highest degree terms if the degree on the top > degree on the bottom, limit DNE if the degree on the top < degree on the bottom, limit is zero.

OpenStudy (anonymous):

ok i'm starting to recall now. thank you both for your help.

OpenStudy (anonymous):

think about if x was a power of ten, say \(x=1000\) then in your example you would have \[\frac{3\times 1000^2+1}{1000}\] which is \[\frac{3000000+1}{1000}\approx3000\] already

OpenStudy (anonymous):

whereas \[\frac{6\times 1000+1}{3\times 1000+8}=\frac{6001}{3001}\approx 2\]

OpenStudy (anonymous):

ok thank you @satellite73. great explanation

OpenStudy (anonymous):

my pleasure